OpenStudy (anonymous):

Find parametric equations for the line through the point that is perpendicular to the given line, and intersects this line. (0, 3, 1) x = 3 + t, y = 3 - t, z = 4t x = -55s y = ? z= ?

OpenStudy (anonymous):

let P(0, 3, 1) and Q(3 + t,3 -t,4t)---> any point where the reqd line intersects the given line

OpenStudy (anonymous):

now drs of PQ is 3+t,-t,4t-1 since PQ is perpendicular to the given line we have (3+t)*1 +(-1)*(-t) +4(4t-1)=0 solving this we get the value of t and hence Q our required line is PQ whose equation can be easily obtained

OpenStudy (anonymous):

here t=1/18 hence x=55s y=3-s z=1-14s